# How to verify that a given process is a Lévy-Driven Ornstein-Uhlenbeck Process

### v1

@@ 4.1 Data and model
We apply our defined verification process to minute-by-minute stock prices of the S&P 500 index constituents from December 2014 until December 2015; the R code for this application can be found in Section 3 of the supplemental material. The S&P 500 is a major stock market index tracking the performance of 500 of the largest publicly traded companies in the United States. It is the benchmark index for large U.S. corporations. Our matrix of data for any given S&P 500 constituent on any given day includes the date given as an integer, the time (minute-by-minute) on a 24 hour clock, and the open price, high price, low, price, and close price from that minute [9].
In order to implement our verification process, we define price spread for two stocks A and B with prices $S_{A}(t)$ and $S_{B}(t)$ as
@@
Spread measures the relative change in the prices of two stocks over time. For modeling the spread dynamics, we refer to the OU process driven by Lévy noise in Equation (2.2). In our example, we calculate the spread dynamics for several large-cap S&P 500 stock constituent pairs in corresponding GICS sectors.
The model of the spread dynamic for the pair Abbot Laboratories (ABT) and Danaher Corporation (DHR) is given in Figure 1:
@@
The model of the spread dynamic for the pair Amgen and Pfizer is given in Figure 2:
@@ 4.2 Verification
We follow our methodology detailed in Section 3. Following Step 1 (see Subsection 3.1), we choose our estimator for $a$. We note that price spread takes on negative values, so we use the LSB estimator to estimate $a$. Following Step 2 (see Subsection 3.2), we choose values for $N$ and $M$. The time period is given to us from the data, namely the observed spread dynamic from December 2014 - December 2015. Thus, we set $M=500$ such that the spread dynamic is sampled at $1/M,2/M,...,N$. Following Step 3 (see Subsection 3.3), we recover the spread dynamic with our chosen $M$ and our estimated $a$ values. Following Step 4 (see Subsection 3.4), we calculate the test statistic and test the correlation of our recovered increments. The application of this methodology on our data can be found in Section 3 of the supplemental material.
We model the spread of several combinations of large-cap stocks in corresponding Global Industry Classification Standard sectors to verify that they are Lévy-driven, having test statistics below the critical value $z_{1-\alpha/2}=1.96$.
We observe varying and indefinite results from our test. An example measuring the price spread between Abbot Laboratories (ABT) and Danaher Corporation (DHR) fails to reject the null hypothesis of uncorrelated increments, producing a test statistic of 1.05. In alignment with this, the spread of Apple (AAPL) and Google (GOOG) fails to reject with a test statistic of 1.85, which suggests that a Lévy driven OU model may be a good model for these spreads. In Figures 3 and 4 below, we plot the recovered increments for the spread of ABT and DHR as a time series and as residuals respectively.
@@
We observe that the recovered increments of the ABT and DHR spread dynamic exhibit relatively stable, uncorrelated behavior, suggesting that the Lévy-driven CAR(1) model could be a good fit. And when plotted as points, the residuals of the recovered increments appear to lack significant trends, outliers, or unusual clustering, further supporting the hypothesis of uncorrelated increments for this spread dynamic.
In Figure 5 below, we plot the sample autocorrelation function for the recovered ABT and DHR spread dynamic.
@@
With a test statistic of 1.05, which is well below the critical value, $z=1.96$, we fail to reject the null hypothesis, which suggest that a Lévy-driven CAR(1) model could be a good fit for this spread dynamic. This is confirmed by the autocorrelation function being well below the dashed line at lag $=$ 1, so we move to Step 5.
Following Step 5 (see Subsection 3.5), we applied our test to assess whether the driving process of these spreads was Brownian motion or not. Following Procedure 1, we found that the spread dynamic of ABT and DHR failed to reject the normality assumption of the recovered increments, suggesting that the increments are consistent with the Brownian motion driven process. We observe consistent results when we apply Steps 1-5 to the spread of Apple (AAPL) and Google (GOOG). Moreover, we apply Procedure 2 to the spread of ABT and DHR testing for Brownian motion, and observe that we spread dynamic still fails to reject the normality assumption under this procedure. Our applications of Procedure 1 and 2 to our data can be accessed in Sections 3.8.1 and 3.8.2 of the supplemental material.
On the other hand, the spread of Goldman Sachs Group (GS) and Morgan Stanley (MS) reject the null, producing a test statistic of 4.30; moreover, the spread of Amgen Inc (AMGN) and Pfizer (PFE) rejected with a test statistic of 4.78, which suggests that a Lévy driven OU model may not be a good model for these spreads. In Figures 6 and 7 below, we plot the recovered increments for the spread of Amgen and Pfizer as a time series and as residuals, respectively.
@@
We observe that there is a very clear leftover relationship in the recovered increments for the Amgen and Pfizer spread dynamic, which suggesting that they are correlated; that suggestion is also supported by the autocorrelation function in Figure 8. The residuals plot reflects clear trends and clustering, indicating potential correlations between increments that do not align with the uncorrelated assumption of a Lévy-driven CAR(1) process.
In Figure 8 below, we plot the sample autocorrelation function for the recovered Amgen and Pfizer spread dynamic.
@@
With a test statistic of 4.78, which is well above the critical value, $z=1.96$, we reject the null hypothesis, which suggest that a Lévy-driven CAR(1) model could be a good fit for this spread dynamic. This is confirmed by the autocorrelation function being well above the dashed line at lag $=$ 1, and we do not need to move on to Step 5.
Results vary largely by GICS sector, but the verification process shows that the Lévy-driven CAR(1) model may not be a good fit for every spread process, including some individual pair combinations as well as several GICS sectors. However, many individual pair combinations as well as several GICS sectors appear to perform well under the verification process, justifying the selection of a Lévy-driven model for their spread processes. Refer to Section 3 of the supplemental material for R code of our spread calculation and use of the verification process and methodology on our data.
5 Example with currency exchange data
@@ 5.1 Data and model
We go on to apply our verification process to realized volatility calculated from Euro/USD currency exchange rate data. This data was taken from Olsen Data, ranging 10 years from June 2007 through June 2017, and contains the Euro/USD bid and ask prices recorded at 5-minute intervals. In the Foreign Exchange Market, the bid price is what a currency dealer is willing to pay for a currency, while the ask price is the rate at which a dealer will sell that currency. We clean out missing data, weekends, fixed holidays, and similar calendar effects, and we are left with 2,463 full days of 5-minute interval data–about 709,555 data entries.
With our cleaned data set, we calculate daily returns, find the corresponding realized variance, and then compute the square root to find daily realized stochastic volatility.
@@
where $r_{t}$ is the daily return at time $t$. Here, realized volatility measures the variation in returns for the Euro/USD currency exchange rate.
Below in Figure 9, we model the Euro/USD returns at each given 5-minute interval which we calculated from our data:
@@
We then follow Equation 5.1 in order to calculate realized volatility. We model the daily realized volatility below in Figure 10:
@@ 5.2 Verification
We then follow our methodology detailed in Section 3 to verify whether the realized volatility is a Lévy-driven CAR(1) process. Following Step 1, we choose the DMB estimator to be our estimator for $a$, as realized volatility only takes on positive values and exhibits characteristics consistent with the Gamma driven and Inverse Gaussian driven processes. Following Step 2, we choose the number of large periods to be $N=35$, and the sampling frequency to be $M=70$, where $2463/M=N$; 2463 being the total number of days in our cleaned data set. Following Step 3, we recover the increments for the Euro/USD realized volatility with our chosen values and estimated parameters. Then following Step 4, we calculate the test statistic and test the correlation of the recovered increments.
@@
We observe that the recovered increments of the Euro/USD realized volatility exhibit noticeable correlation over the time period, suggesting that the increments are not independent. When plotted as individual points, the residuals show a noticeable pattern indicating that the increments may exhibit correlation, which suggest that a Lévy-driven CAR(1) model may not be a good fit; that suggestion is supported by the autcorrelation function in Figure 13.
In Figure 13 below, we plot the sample autocorrelation function for the recovered Euro/USD realized volatility.
@@
With a test statistic of 3.76 for the Euro/USD realized volatility, which is well above the critical value, $z=1.96$, we reject the null hypothesis, which suggest that a Lévy-driven CAR(1) model could be a good fit for this spread dynamic. This is confirmed by the autocorrelation function being well above the dashed line at lag $=$ 1, and we do not need to move on to Step 5.
As the test rejects the Lévy-driven CAR(1) model, we do not need to go on to Step 5. However, in alignment with [5], we believe that the Lévy-driven CARMA(2,1) process could be a good fit for realized volatility.
@@ 6 Conclusion
In conclusion, our verification process proves very useful in testing the fit of models assumed to be Lévy-driven. In this paper, we test a Lévy-driven model for price spread between S&P 500 stock pairs, as well as Euro/USD exchange data, but the verification process can be applied to other real world data, e.g., interest rates, credit risk, population dynamics, and more. Our test yields strong results, failing to reject the null hypothesis of uncorrelated increments in most cases for many simulated Lévy-driven CAR(1) background processes, i.e., Brownian motion, Gamma, Beta, Inverse Gaussian, and mixed combinations of these processes. Moreover, we observed that when testing for Brownian motion, Procedure 1 and Procedure 2 rejected at high rates when the recovered increments were different from the normal distribution for varying combinations of $N$, $M$, and $a$, and the test rejected at p-values near 0.05 under the $H_{0}$. Our methodology proves a powerful tool in verifying whether a given process is an Ornstein-Uhlenbeck process and has direct application to real world data. This paper details the precise methodology to apply this test, and our supplemental material includes further explanation as well as R code that corresponds with the verification process methodology and our example with financial data. Going forward in our future work, we plan to write a similar, user accessible, methodological paper detailing how to verify that a given process is a Lévy-driven CARMA(2,1) process.

### v2

@@ 4.1 Data and model
We apply our verification process to minute-by-minute stock prices of the S&P 500 index constituents from December 2014 to December 2015. The R code for this application is provided in Section 3 of the supplemental material. The S&P 500 is a major stock market index that tracks the performance of 500 of the largest publicly traded companies in the United States, serving as the benchmark index for large U.S. corporations. For each S&P 500 constituent on any given day, our data matrix includes the date (represented as an integer), the time (minute-by-minute on a 24-hour clock), and the open, high, low, and close prices for each minute [9].
To implement our verification process, we define the price spread between two stocks, A and B, with prices $S_{A}(t)$ and $S_{B}(t)$ as follows:
@@
Spread measures the relative change in the prices of two stocks over time. The goal is to evaluate whether these spread dynamics can be modeled using a Lévy-driven CAR(1) process, Equation (2.2), as outlined in our methodology. In our example, We calculate the spread dynamics for several large-cap S&P 500 constituent pairs within their corresponding Global Industry Classification Standard (GICS) sector (GICS) sectors.
The model for the spread dynamics of the pair Abbott Laboratories (ABT) and Danaher Corporation (DHR) is shown in Figure 1:
@@
The model for the spread dynamics of the pair Amgen and Pfizer is shown in Figure 2:
@@ 4.2 Verification
We follow the methodology detailed in Section 3. In Step 1 (see Subsection 3.1), we select our estimator for $a$. Since the price spread can take on negative values, we use the LSB estimator to estimate $a$. In Step 2 (see Subsection 3.2), we determine the values for $N$ and $M$. The time period is provided by the data, specifically the observed spread dynamics from December 2014 to December 2015. Therefore, we set $M=500$ so that the spread dynamics are sampled at intervals of $1/M,2/M,...,N$. In Step 3 (see Subsection 3.3), we recover the spread dynamics using our chosen $M$ and estimated $a$. In Step 4 (see Subsection 3.4), we calculate the test statistic and test the correlation of the recovered increments. The application of this methodology to our data is presented in Section 3 of the supplemental material.
We model the spread of several combinations of large-cap stocks from corresponding Global Industry Classification Standard (GICS) sectors to verify that they are Lévy-driven, with test statistics falling below the critical value. $z_{1-\alpha/2}=1.96$.
We observe varying and sometimes inconclusive results from our test. For example, when measuring the price spread between Abbott Laboratories (ABT) and Danaher Corporation (DHR), we fail to reject the null hypothesis of uncorrelated increments, with a test statistic of 1.05. Similarly, the spread between Apple (AAPL) and Google (GOOG) also fails to reject the null hypothesis, yielding a test statistic of 1.85. This suggests that a Lévy-driven Ornstein-Uhlenbeck (OU) model may be a suitable model for these spreads. In Figures 3 and 4 below, we present the recovered increments for the ABT-DHR spread both as a time series and as residuals, respectively.
@@
We observe that the recovered increments of the ABT-DHR spread dynamics exhibit relatively random, uncorrelated behavior, suggesting that the Lévy-driven CAR(1) model is a good fit. When plotted as points, the residuals of the recovered increments show no significant trends, outliers, or unusual clustering, further supporting the hypothesis of uncorrelated increments for this spread dynamic.
In Figure 5 below, we plot the sample autocorrelation function (ACF) for the recovered ABT-DHR spread dynamics.
@@
With a test statistic of 1.05, well below the critical value of $z=1.96$, we fail to reject the null hypothesis, suggesting that a Lévy-driven CAR(1) model could be a good fit for this spread dynamic. This is further supported by the autocorrelation function, which remains well below the dashed line at lag $=$ 1, allowing us to proceed to Step 5.
Following Step 5 (see Subsection 3.5), we applied our test to assess whether the driving process of these spreads was Brownian motion. Using Procedure 1, we found that the spread dynamic of ABT and DHR failed to reject the normality assumption of the recovered increments, suggesting that the increments are consistent with a Brownian motion-driven process. We observed consistent results when applying Steps 1-5 to the spread of Apple (AAPL) and Google (GOOG). Additionally, for further confirmation, when we applied Procedure 2 to the spread of ABT and DHR to test for Brownian motion, the spread dynamic still failed to reject the normality assumption. Our applications of Procedures 1 and 2 to the data can be found in Sections 3.8.1 and 3.8.2 of the supplemental material.
On the other hand, the spread between Goldman Sachs Group (GS) and Morgan Stanley (MS) rejects the null hypothesis, producing a test statistic of 4.30. Similarly, the spread between Amgen Inc. (AMGN) and Pfizer (PFE) also rejects the null hypothesis with a test statistic of 4.78. This suggests that a Lévy-driven OU model may not be a good fit for these spreads. In Figures 6 and 7 below, we plot the recovered increments for the spread between Amgen and Pfizer, both as a time series and as residuals, respectively.
@@
We observe a very clear leftover relationship in the recovered increments for the Amgen and Pfizer spread dynamic, suggesting that they are correlated. This observation is further supported by the autocorrelation function in Figure 8. The residuals plot shows clear trends and clustering, indicating potential correlations between increments that do not align with the uncorrelated assumption of a Lévy-driven CAR(1) process.
@@
With a test statistic of 4.78, which is well above the critical value of $z=1.96$, we reject the null hypothesis, suggesting that a Lévy-driven CAR(1) model may not be a good fit for this spread dynamic. This conclusion is further supported by the autocorrelation function, which is well above the dashed line at lag $=$ 1. Therefore, we do not need to proceed to Step 5.
Conclusion:
The results vary significantly by GICS sector, but the verification process indicates that the Lévy-driven CAR(1) model is not be a good fit for all spread processes. This includes certain individual pair combinations as well as some GICS sectors. However, many individual pair combinations and several GICS sectors perform well under the verification process, supporting the selection of a Lévy-driven model for their spread processes. For further details, including R code for spread calculation and the application of the verification methodology to our data, refer to Section 3 of the supplemental material.
5 Application to Realized Volatility of Euro/USD Exchange Rate
@@ 5.1 Data and model
We next apply our verification process to the realized volatility calculated from Euro/USD currency exchange rate data. This data, sourced from Olsen Data, spans 10 years (from June 2007 to June 2017) and consists of Euro/USD bid and ask prices recorded at 5-minute intervals. In the Foreign Exchange (FX) market, the bid price represents the amount a currency dealer is willing to pay for a currency, while the ask price is the rate at which the dealer is willing to sell it.
After cleaning the data to remove missing entries, weekends, fixed holidays, and other calendar effects, we are left with 2,463 full trading days of 5-minute interval data, totaling approximately 709,555 data points.
With our cleaned dataset, we calculate daily returns, determine the corresponding realized variance, and then compute the square root to obtain the daily realized stochastic volatility.
@@
where $r_{t}$ is the daily return at time. In this context, realized volatility measures the variation in returns for the Euro/USD currency exchange rate. Below, in Figure 9, we model the Euro/USD returns for each 5-minute interval, calculated from our data:
@@
We then follow Equation (5.1) to calculate realized volatility. The daily realized volatility is modeled below in Figure 10:
@@ 5.2 Verification
We then follow the methodology outlined in Section 3 to verify whether the realized volatility follows a Lévy-driven CAR(1) process. In Step 1, we choose the DMB estimator for $a$, as realized volatility only takes on positive values and exhibits characteristics consistent with Gamma-driven and Inverse Gaussian-driven processes. In Step 2, we set the number of large periods to $N=35$ and the sampling frequency to $M=70$, where $\frac{2463}{M}=N$, with 2463 representing the total number of days in our cleaned dataset. In Step 3, we recover the increments for the Euro/USD realized volatility using the chosen values and estimated parameters. Finally, in Step 4, we calculate the test statistic and test the correlation of the recovered increments.
@@
We observe that the recovered increments of the Euro/USD realized volatility exhibit noticeable correlation over time, suggesting that the increments are not independent. When plotted as individual points, the residuals show a distinct pattern, indicating that the increments may exhibit correlation. This suggests that a Lévy-driven CAR(1) model may not be a good fit, a conclusion supported by the autocorrelation function in Figure 13.
With a test statistic of 3.76 for the Euro/USD realized volatility, which is well above the critical value $z=1.96$, we reject the null hypothesis. This suggests that a Lévy–driven CAR(1) model may not be a good fit for this spread dynamic. This is confirmed by the autocorrelation function being well above the dashed line at lag $=$ 1, and thus, we do not need to proceed to Step 5.
@@
In line with [5], we believe that the Lévy-driven CARMA(2,1) process could be a good fit for realized volatility. However, further verification is needed to confirm that the Lévy-driven CARMA(2,1) model is indeed a suitable fit, and this work will be conducted in the near future.
@@ 6 Conclusion
In conclusion, our verification process proves highly effective in testing the fit of models assumed to be Lévy-driven. In this paper, we apply the process to test a Lévy-driven model for the price spread between S&P 500 stock pairs and Euro/USD exchange rate data. However, the verification process can be extended to other real-world data, such as interest rates, credit risk, population dynamics, and more.
Our tests yield strong results, failing to reject the null hypothesis of uncorrelated increments in most cases, approximately 95% of the time at the nominal level ($\alpha=0.05$), across many simulated Lévy-driven CAR(1) background processes, including Brownian motion, Gamma, Beta, Inverse Gaussian, and mixed combinations of these processes. Additionally, when testing for Brownian motion, Procedures 1 and 2 produced p-values approaching 0.05 under $H_{0}$. Moreover, when testing the power of these two procedures, they rejected the normal distribution hypothesis at high rejection rates when the recovered increments deviated from a non-normal distribution for varying combinations of $N$, $M$, and $a$.
Our methodology proves to be a powerful tool for verifying whether a given process is an Ornstein-Uhlenbeck process, with direct application to real-world data. This paper provides the precise methodology for applying this test, and our supplementary material includes further explanations as well as R code corresponding to the verification process and the data examples.
In future work, we plan to write a user-accessible methodological paper detailing how to verify that a given process is a Lévy-driven CARMA(2,1) process.